# Crossroads of Continents: Automated Artifact Extraction for Cultural Adaptation with Large Multimodal Models

### v1

@@ 4.1 Methods
We use the open vocabulary object detection Zareian et al. (2021) capabilities of GPT-4V vision-preview, formulating a detailed prompt (see Appendix C.2) to extract information about the target concept class in a Dalle Street image, including text descriptions, color, [^5] Note that this corresponds to appearance of different objects and persons in images, and does not imply race. and person count. GPT-4V’s strong instruction following capabilities allow the generation of a nearly perfectly validated JSON string, which we lightly post-process, and then provide that as input to GPT-4 turbo to get a summary of the objects for any given image. In our initial experiments, GPT-4V outperformed LLaVA significantly in terms of following complex instructions for object detection, so we only report results from GPT-4V.
Figure 5: We explore how countries are distributed on a color spectrum by first calculating a global average RGB vector and then defining deltas along each axes aggregated at the country level. Takeaway: We find interesting associations - Greece is strongly correlated with blue, Burkina Faso with red.
This process results in many unique cultural artifacts associated with a single country. To narrow this down to the salient artifacts that occur relatively more frequently for one country but less for others (i.e. stereotypical associations), we follow an approach similar to Jha et al. (2024): we compute the term frequencies of each artifact for each country and also compute document frequency as the number of times an artifact occurs across all countries. The final tf-idf score for a given artifact for a particular country is obtained by multiplying term frequency and the inverse of the document frequency. We then use the distribution of these scores by calculating the mean and standard deviation across all scores, and then perform a qualitative evaluation of both outliers and scores that fit the global distribution.
Figure 6: We identify more than 18,000 unique cultural artifacts across all countries as part of our second task, and then filter them to find salient ones. This figure shows strongest correlated artifacts for 5 randomly picked countries. More examples in Appendix C.4.13.
Color Associations for Countries
We calculate the mean RGB vector for each synthetic Dalle Street image, and then calculate the average to define a global mean vector. We repeat this process at the country level to obtain a mean vector for each country. Then, we find the distance of each country from the global mean and split across the three dimensions of the vector to find colors that are more likely to be associated with some countries. One of the challenges of this approach is that we only consider a limited number of color axes which may not be sufficient to capture finer signals about more complicated color associations, similar to McCarthy et al. (2019). We leave this for future work.
Counting the Number of People
In our initial observations, we noticed that DALL-E 3 tends to generate different population densities in images from different countries for otherwise identical prompts. To explore this, our object detection prompt also reports the count of people for each image, split in two buckets (between $1$ and $10$ people, or more than $10$ people). We further divide the first bucket into two smaller buckets based on the exact count (less than $5$ people, or between $5$ and $10$ people), to get $3$ buckets overall. We process this output to consider all possible variations of these terms (for example, people, person, man, woman, etc) and aggregate statistics at the country level to find the distribution of population densities in generated images.
@@ 4.2 Results
We analyse all the DALL-E 3 generated images using GPT-4V to discover hidden associations that these LMMs makes between countries and cultural artifacts.
Selected Examples of Associations within Standard Deviation
Figure 6 (and Figure C.4.13) illustrates some of the randomly sampled interesting cultural artifacts we discover, and Figure 4 shows the distribution of the tf-idf scores we calculate for each country-artifact pair. Artifacts that fall outside the range as denoted by the standard deviation lines from the mean, would be very strongly associated with a particular country and are thus very important to more closely look at, to consider if they are stereotypical (especially negatively) or not. We include statistics for the number of artifacts we identify along with more examples of interesting associations in Appendix C.4 Table LABEL:tab:artifact-interesting.
Color associations
We find that models tend to not only associate particular cultural artifacts with countries, but also colors. Figure 5 (and Figure C.4.18) shows a subset of all the countries in the Dalle Street images versus the delta values in the RGB vector components from the global average vector. Once again, for some countries, these delta values fall well outside the range denoted by the standard deviation lines from the mean and these represent strong presence of that color component in images for that country. For example, Greece is very strongly linked to the colors blue and green, whereas Indonesia and Netherlands are associated with red. An explanation for why this may be happening may be inferred from inspecting the generated images. Images of Greece typically feature a lot of sea, blue decorations, and blue patterns (usually stereotypical) whereas Netherlands is famous for its red tulip fields, which is also an artifact association that we are able to obtain from our previous analysis.
Figure 8: Our CultureAdapt pipeline identifies the fine-grained elements of an image that should be modified for cultural adaptation. Note that any in-painting technique can be used. Compare CultureAdapt with DALL-E 3, which generates a completely different image. More examples of edits in Appendix C.4.14.
People-Count Associations
We parse the information obtained about population density per DALL-E 3 image from our object detection prompt for GPT-4V, to get $3$ buckets overall (between $1$ and $5$ people, between $5$ and $10$ people or more than $10$ people). Figure 7 (and  C.4.19) shows a sample of countries for each bucket versus the number of images for that country which fall into that bucket.
A general trend is that images usually fall into the two extreme buckets, and we have very few occurrences of the middle bucket. Cameroon gets the highest number of occurrences in the bucket of more than $10$ people counts, despite other countries like China and India having much higher population densities in the real world. In general, African countries fall into the high-person-count buckets, while European ones fall largely in the few-people one. This might reflect the models’ perception of societies as largely individualistic or collectivist. That said, as there might be errors by GPT-4V in the person-counting process in the images –which we have not validated except from looking at a handful of examples– we recommend taking these results with a grain of salt, and leave this further exploration for future work.
5 Cultural Adaptation (Task 3)
We propose a highly composable method, CultureAdapt, for fine-grained editing that aims to use the discovered cultural artifacts from Task $2$, and adapts a given source image to a target culture. This has many potential downstream applications. For example, in generative AI applications, if the generated content is detected to not follow the specified country label, then our pipeline can be used to update it in near real-time. Another useful application of our pipeline would be as a source of data augmentation for pre-training corpora. We can use the pipeline to create culturally augmented data and also include translations to create a parallel corpus across languages and cultures which can then be used for continued pre-training to mitigate identified issues.
5.1 Methods: CultureAdapt
Recent work on cultural translation Khanuja et al. (2024); Li and Zhang (2023); Fung et al. (2024) defines different approaches for adapting images or text from one culture to another. We propose a simple automated and highly composable pipeline that can extract culture specific entities from an image, find the most salient ones, ground them with bounding boxes, and then modify only these objects without affecting the rest of the image. After this cultural adaptation process, we use CLIPScore Hessel et al. (2021) as a reference-free metric to measure the image-country compatibility using CLIP Radford et al. (2021) (treating the name of the country as the caption).
Our pipeline is built with fully customizable components. For example, we currently use GPT-4V vision-preview for open-vocab object detection, Grounding DINO Liu et al. (2023b) for grounding objects in bounding boxes and Stable Diffusion $2$ inpainting Rombach et al. (2021) for updating these grounded object masks. But, we could also use something like Tag2Text Huang et al. (2023) or RAM Zhang et al. (2023) to generate captions for an image, and then extract tags from these captions corresponding to objects, which can then be passed into Grounded SAM Ren et al. (2024) to obtain bounding boxes and segmentation masks, which will finally be used by any inpainting model like Stable Diffusion 3 AI (2024) or MimicBrush Chen et al. (2024). We demonstrate the feasibility of our approach.
Evaluation
Let image $I_{1}$ originally correspond to country $C_{1}$. We use CultureAdapt to translate it to $I_{2}$ for country $C_{2}$. The CLIPScore of an image-country pair is denoted as $S(I,C)$. We define two delta values from this as follows:
$\Delta_{1}$ $=S(I_{2},C_{1})-S(I_{1},C_{1})$
$\Delta_{2}$ $=S(I_{2},C_{2})-S(I_{1},C_{2})$
If $\Delta_{1}<0$ and $\Delta_{2}>0$, then it would imply that after cultural adaptation, the image is closer to the target country $C_{2}$ than the source country $C_{1}$. This is exactly the success criterion for our task.
5.2 Results
We show examples of using CultureAdapt in Figure 8, where we start with a Dalle Street image for the concept of front door in the country Greece. Then we culturally translate these to $3$ other countries, China, India and USA. For all of these images shown, $\Delta_{1}<0$ and $\Delta_{2}>0$ (values in Appendix C.4.3), implying that quantitatively we have measured a culture change for all 3 cases. We include more examples of CultureAdapt in Appendix C.4.14.
We further contrast these with images generated for the same concept, country pairs from DALL-E 3 to highlight differences in the two approaches. Images generated from DALL-E 3 may be more representative of the target country as per CLIPScore but they are stylistically very different from the original source image (i.e lack object consistency before and after the cultural translation process), thus making them less usable for different applications where this property is necessary.
6 Human Studies
We choose a small subset of our data, about $300$ images randomly sampled, and perform a couple of human studies with $14$ people. Here we present highlight findings due to space, with more details and annotator demographics in Appendix C.3.
Country Representation in Dalle Street
We examine whether annotators consider specific images to be stereotypical representations of a country. Figure C.4.20(a) shows that most participants agree or strongly agree, with only $2$ out of $304$ images receiving disagreement. When unsure, participants tend to neither agree nor disagree about the appropriateness of an image, . This study indicates that our generated images effectively represent country-level associations. Our participants also mark this information with visual cues (Table C.LABEL:tab:artifact-human), from explicit (e.g. flags) to implicit (e.g. color choice, nuanced cultural artifacts). Based on this analysis, we develop our framework for cultural artifact extraction and cultural adaptation using Dalle Street images.
Human baseline for Cultural Awareness
We look at accuracy metrics from $5$ different perspectives. For our study, we allowed users to select a minimum of $1$ label and a maximum of $5$ per image. These labels can be at the country level, or at the continent level or subcontinent level. So, we can evaluate exact match accuracy first at the country level, then at the subregion level and finally at the continent level. In addition, we also consider the case of union (correct answer at one of the three levels is present in the human provided labels), and the case of intersection (correct answer at all three levels are present in the human provided labels). Figure C.4.20(b) shows that on average participants have a low accuracy at the country level, but nearly double performance every time we go up a geographical level, which is along the lines of what we expect.
@@ 7 Conclusion
This study addresses the critical need for cultural awareness in Large Multimodal Models (LMMs) by introducing a comprehensive framework to evaluate and enhance their cultural competence. We create a large-scale, culturally diverse dataset of $9,935$ images across $67$ countries and $10$ concept classes, facilitating robust evaluation of LMMs on cultural awareness tasks. Further, we introduce a cultural artifact extraction task to identify over $18,000$ cultural artifacts that co-occur frequently with these countries, revealing significant insights into the implicit cultural associations encoded in these models. We also propose a cultural adaptation task and a highly composable pipeline CultureAdapt to adapt images across cultural contexts with fine-grained edits. Overall, this work emphasizes the importance of developing culturally sensitive AI systems and provides a foundational benchmark for future research towards improvement in cultural representation.
@@
Coverage
Our work focuses on a limited number of concept classes across a small portion of all countries in the world. While we are able to cover most geographic regions and sub-regions with this data, it is important to ensure wider coverage over time to ensure that no culture is left behind.
Open Vocabulary Object Detector
Our pipeline for automated cultural adaptation is only as good as its components. The first and most crucial component of meaningful object tag extraction is approached using GPT-4V in our work, but open-source alternatives like RAM and Tag2Text are slowly catching up to cover more wider variety of classes.
Stable Diffusion Biases
While we are using Stable Diffusion for inpainting masked objects only, our approach does not guarantee that the object generated from the diffusion process will not reflect any harmful stereotypes. Our CLIPScore based metric for measuring country deltas is also preliminary and limited by the strength of the underlying CLIP model embedddings and its inherent biases as well.
Mitigation Strategies
An important aspect of studying differences in culture awareness of LMMs is understanding how to develop approaches to mitigate them. This may involve looking more closely at pre-training corpora and developing efficient techniques for data augmentation for the same, so that secondary approaches like prompting in local languages can extract more aligned responses. While exploring all of these is beyond the current scope of our paper, we present a survey of the most relevant recent work that attempts to do so.
Attributing model behavior to specific training examples
Methods like Park et al. (2023) aim to better understand issues with training data by assigning scores to data points which cause the model to behave in a certain way. This is useful as a first step to trace training data that reflects potentially harmful stereotypes from the list of automatically identified cultural artifacts that we identify. Recent work Hall et al. (2023) has looked at the LAION corpus Schuhmann et al. (2022) to determine that geographical regions may only co-occur with more stereotypical image captions. For example, a Greek person having dinner in Greece is less likely to caption a photo of that scene as “having food in Greece” but a tourist is more likely to mention something like “when in Greece, you must moussaka” in a caption for the same image, leading to perpetuation of stereotypes when these image-caption pairs are used for training models. Similarly, color associations that we find in this paper may simply be because images were scraped from travel websites which usually portray a more vibrant view of a country than the native viewpoint.
Data Augmentation
There are many interesting approaches to generate higher quality data to mitigate cultural differences. Dollar Street for example uses photographers as a source of collecting this data manually, whereas MaRVL uses crowdsourcing. Other automated approaches include knowledge acquisition by scraping from internet sources Fung et al. (2024), simulating conversations between LLMs Li et al. (2024b), semantic data augmentation Li et al. (2024a) and value preference aligment by collecting preferences during conversations with LLMs Kirk et al. (2024). For multimodal data augmentation, methods include approaches based on finding synonyms from semantic graphs of image captions Li and Zhang (2023) and end-to-end pipelines for image transcreation Khanuja et al. (2024). Our approach CultureAdapt aims to contribute to this area of research by automatically identifying cultural artifacts and providing a composable pipeline for finer-grained edits to these artifacts.
Prompting in local languages
Hall et al. (2023) find that images generated with non-English languages tend to struggle with prompt consistency and continue to show stereotyped representations, whereas Pouget et al. (2024) find that pre-training with global, unfiltered data before fine-tuning on English content can improve cultural understanding without sacrificing performance on popular benchmarks. These contrasting findings imply the need to dedicate more efforts to find the impact of pre-training on data that has been augmented for visual modality using pipelines such as ours along with translated image captions for a multimodal parallel corpora Etxaniz et al. (2024) to eliminate geographical stereotypes. Improving the quality of the text encoders used in LMMs is another important angle to look at in this light.
@@ Acknowledgements
Appendix A Related work
Recent spike in interest for culturally aware NLP has produced many relevant works, which has inspired different aspects of our research.
Li and Zhang (2023) uses semantic graphs to develop an annotation free data augmentation approach for augmenting cultural components of captions. The main challenge of this approach is that the cultural artifacts identified for replacement are copy-pasted into the target image, which leads to inconsistencies at the object boundaries. Similarly, Khanuja et al. (2024) define the image transcreation task formally in the context of culture, and develop pipelines for cultural translation. The edited images from these pipelines tend to vary in structure, and may sometimes look very different from the source image. Our proposed CultureAdapt pipeline overcomes both of these challenges, by using automatically generated object labels grounded with bounding boxes to create precise semantic masks which are then filled in with a diffusion based model in an inpainting setting.
Qiu et al. (2024) calculates co-occurence statistiscs for features extracted from images and Jha et al. (2024) assigns importance scores to attributes to determine those which co-occur more frequently for an identity group over others. We explore both of these ideas in our work to understand which cultural artifacts are more likely to co-occur for a given country. Pouget et al. (2024) and Hall et al. (2023) both explore the two real-world datasets, MaRVL and Dollar Street, specifically developing reliable metrics to measure geo-localization and object consistency across regions. Liu et al. (2024) defines a taxonomy for culturally aware and adapted NLP which we borrow terminology from, whereas Adilazuarda et al. (2024) provides a survey of the current state of research in this sub-field and also what is yet to be done.
Other works also explore downstream tasks, for example, Ignat et al. (2024) looks at cross-cultural inspiration detection in social media data, Seaborn et al. (2024) develops voice assistants with cultural sensitivity and Srinivasan et al. (2024) improves cultural diversity of search results.
Appendix B Dalle Street generation prompt
Appendix C Appendix
C.1 Dataset details

### v2

@@ 5.1 Methods
Recent works on cultural translation Khanuja et al. (2024); Li and Zhang (2023); Fung et al. (2024) define different approaches for adapting images or text from one culture to another.
CultureAdapt
Our pipeline (Figure 10) uses GPT-4V for open-vocab object detection to extract implicit cultural associations from the source image. Next, we use Grounding DINO Liu et al. (2023b) to ground these objects with bounding boxes, which we convert into masks. We then create an inpainting prompt by adding the target country to the list of detected objects and use Stable Diffusion $2$ inpainting Rombach et al. (2021) to edit [^4] We acknowledge that the resulting image may represent common stereotypes because the underlying artifacts may be implicitly associated with a stereotypical view of the country. the image by filling in the masked pixels. We evaluate our method using CLIPScore Hessel et al. (2021) to measure image-country similarity (treating the name of the country as the caption) and cosine similarity of DINO-ViT Caron et al. (2021) embeddings for measuring structural preservation. Our choice of metrics is inspired from  Khanuja et al. (2024).
Modularity
Our pipeline is modular, allowing components to be easily swapped to improve performance over time. For example, we could use Tag2Text Huang et al. (2023) or RAM Zhang et al. (2023) for image captioning, extract object tags, and then use Grounded SAM Ren et al. (2024) for bounding boxes and segmentation masks. These can then be passed to inpainting models like Stable Diffusion 3 AI (2024) or MimicBrush Chen et al. (2024).
Figure 10: The CultureAdapt pipeline first identifies objects in an image using GPT-4V and grounds them with bounding boxes from Grounding DINO. These masks, along with a prompt containing object names and the target country, are used by Stable Diffusion for inpainting and cultural adaptation. Takeaway: Unlike DALL-E 3 (which generates a completely new image) or other editing methods Khanuja et al. (2024) that struggle with cultural adaptation, CultureAdapt makes precise, meaningful edits. Additional examples in the Appendix (Figure 29).
Baseline
The closest related work is by Khanuja et al. (2024), which proposes three methods for image editing. We compare our approach with their two most relevant methods, providing qualitative examples (Figure 10) and quantitative evaluations (Table 7). Additionally, we conduct a qualitative study assessing human preferences for layout preservation and cultural relevance changes (Figure 12) for country pairs common to both studies. We also perform extensive statistical testing to compare structural similarity and editing success, using CLIPScore-based metrics, for both approaches.
Evaluation
Let image $I_{1}$ correspond to country $C_{1}$ and $I_{2}$ be its adaptation for country $C_{2}$. The CLIPScore for an image-country pair is denoted as $S(I,C)$. We define two deltas:
$\Delta_{1}$ $=S(I_{2},C_{1})-S(I_{1},C_{1})$ (1)
$\Delta_{2}$ $=S(I_{2},C_{2})-S(I_{1},C_{2})$ (2)
If $\Delta_{1}<0$ and $\Delta_{2}>0$, it indicates successful adaptation, where $I_{2}$ is closer to $C_{2}$ than $C_{1}$. Our primary metric, $M_{1}$, tracks how often this condition is met. A secondary metric, $M_{2}$, compares $\Delta_{2}-\Delta_{1}$ to evaluate success, though it’s less ideal for cases where $\Delta_{1}$ is positive. We also use another metric, SSIM, to measure structural similarity between $I_{1}$ and $I_{2}$ via DINO-ViT embeddings. A good edit performs well across editing and similarity metrics.
@@ 5.2 Results
Empirical findings indicate that our method works well both qualitatively and quantitatively.
Qualitative comparisons
In Figure 10, we show CultureAdapt applied to a randomly selected image, for adapting from Greece to China, India, and the USA. By visually contrasting with results from DALL-E 3, we see that our approach preserves structural similarity better, and by comparing with the baselines e2e-instruct and cap-edit, we demonstrate our method’s ability to make meaningful edits. Figure 29 includes more examples.
Comparison with baseline
We compare CultureAdapt with cap-edit (Figure 11) for $20$ country pairs (Table 7). Both methods produce images similar to the source, but CultureAdapt edits are overall more culturally relevant.
To test this empirically, we compare similarity scores using the Wilcoxon signed-rank test (Shapiro-Wilk: $p=5.99\times 10^{-39}$, indicating non-normality). cap-edit has a slightly higher mean similarity $0.97$ than CultureAdapt $0.94$, with a statistically significant difference ($p=6.02\times 10^{-215}$) for ($\alpha=0.05$). The bootstrapped $95$% confidence interval for the mean difference is [$0.0298$, $0.0333$], which does not include 00, supporting this result.
For editing metrics, CultureAdapt outperformed cap-edit in a statistically significant way. For $M_{1}$, $54\%$ of samples met the condition versus $50\%$ for cap-edit ($p=7.43\times 10^{-5}$, McNemar’s test). For $M_{2}$, CultureAdapt had a higher mean score ($3.11$ vs. $2.68$), with the difference significant per the Wilcoxon test ($p=4.39\times 10^{-11}$), as non-normality was confirmed by the Shapiro-Wilk test ($p=1.23\times 10^{-16}$). The bootstrapped $95\%$ confidence interval for $M_{2}$ differences supports our conclusion ([$-0.560$, $-0.307$]). In a human study of $100$ images, $3$ participants rated both methods equally preferable on structure preservation and cultural relevance (Figure 12).
Figure 11: Our method performs better than other approaches in terms of editing metrics while still maintaining comparable structural similarities.
Figure 12: When asked to compare editing outputs from CultureAdapt to the existing baseline cap-edit, our annotators have nearly similar preferences for both methods in terms of (1) the ability of the methods to maintain layout similarity and (2) cultural relevance of the edited image to the target country.
Error Analysis
We identify two common error modes (Figure 13): ($1$) multiple similar objects to be edited in the source image lead to masks covering a significant portion of the image, often in an overlapping manner, resulting in major changes as the diffusion process needs to inpaint from scratch, and ($2$) when editing realistic DALL-E 3 images, the edited objects sometimes lose realism. Replacing Grounding DINO with segmentation models (e.g., Grounded SAM) helps mitigate the first issue, but the second remains an open problem, as noted by others Hall et al. (2023). Other less obvious issues include not maintaining the correct count or orientation of objects and not generating human faces correctly, as the underlying diffusion model is trained with a privacy filter.
Figure 13: Common error cases with our method occur when there are multiple objects in the image or when the image is generated in a realistic style with DALL-E 3.
6 Related Work
The growing interest in culturally aware NLP has inspired various aspects of our research.
Li and Zhang (2023) propose a data augmentation approach using semantic graphs to enhance cultural components in captions. However, their method often results in inconsistencies at object boundaries when cultural artifacts are copied and pasted into images. Similarly, Khanuja et al. (2024) formalize the task of image transcreation, but their pipelines can produce images that differ significantly from the source or not at all. Our CultureAdapt pipeline maintains image coherence by generating semantic masks with bounding boxes and using diffusion-based inpainting.
Qiu et al. (2024) calculate co-occurrence statistics from image features, and Jha et al. (2024) assign importance scores to attributes that frequently co-occur for identity groups. We build on these ideas to identify salient cultural artifacts likely to co-occur for a given country. Liu et al. (2024) provide a taxonomy for culturally aware NLP, from which we adopt terminology (cultural artifacts and adaptation). Pouget et al. (2024) and Hall et al. (2023) both explore the two real-world datasets, MaRVL and Dollar Street, specifically developing reliable metrics to measure geo-localization and object consistency across regions.
@@ 7 Conclusion
This study addresses the critical need for cultural awareness in Large Multimodal Models by introducing a comprehensive framework to evaluate and enhance their cultural competence. We create a large-scale, culturally diverse dataset of $9{,}935$ images across $67$ countries and $10$ concept classes, facilitating benchmarking of LMMs on cultural awareness tasks. Further, we introduce an artifact extraction task to identify over $18{,}000$ artifacts that co-occur frequently with these countries, revealing significant insights into the implicit cultural associations encoded in these models. We also propose CultureAdapt, a pipeline to adapt images across cultural contexts with fine-grained edits. Overall, this work emphasizes the importance of developing culturally sensitive AI systems and provides a foundational benchmark for future research toward improvement in cultural representation.
@@
Our experiments use a specific version of a closed-source API model, but later versions usually offer similar or improved performance at lower costs. The effectiveness of our cultural adaptation approach depends on its components, particularly object tag extraction, which currently works well when using GPT-4V. Open-source alternatives like RAM and Tag2Text would become more viable over time for this task.
Our method could be extended for data augmentation to improve cultural awareness in models, though this falls outside the scope of our current work. Additionally, exploring multilingual prompts to analyze associations across languages would be interesting but is not our focus here. Since we use Stable Diffusion for inpainting, we cannot fully prevent the stereotypical generations. These challenges point to the need for more culturally nuanced models, as discussed in recent works Park et al. (2023); Hall et al. (2023); Li et al. (2024).
@@ Acknowledgements
Appendix A Appendix
  1. 1.
    Dataset details A.1
    1. (a)
      Dollar Street concept classes
    2. (b)
      Dollar Street countries
    3. (c)
[…]
    4. (d)
[…]
    5. (e)
      MaRVL Country to Language Mappings
  2. 2.
    Prompt details A.2
    1. (a)
      Dalle Street generation
    2. (b)
      Cultural Awareness classifier
    3. (c)
      Object Detection
  3. 3.
    LLM hyperparameters A.3
    1. (a)
      Generation settings
    2. (b)
      Computation budget
  4. 4.
    Human Study - the Annotators A.4
    1. (a)
      Annotator Demographics
    2. (b)
      How many studies do we have
    3. (c)
      Total number of annotations
  5. 5.
    Human Study - the Interfaces A.5
    1. (a)
      General Instructions for Study 1 and 2
    2. (b)
      Study 1: Verify the quality of generated images
      1. i.
        Task 1 Instructions
    3. (c)
      Study 2: Human performance on our cultural awareness benchmark
      1. i.
        Task 2 Instructions
      2. ii.
        Anonymized performance of individual Human Study participants
    4. (d)
      Study 3 and 4: Artifact hallucination and cultural relevance
    5. (e)
      Study 5: People count associations
    6. (f)
      Study 6 and 7: Layout preservation and cultural relevance after editing
  6. 6.
    Additional Results A.6
    1. (a)
      Task 1 - Cultural awareness
      1. i.
        Performance of LLaVA on income quartiles for Dollar Street
      2. ii.
        Confusion matrices for LLaVA and GPT-4V on Dalle Street images
      3. iii.
        Confusion matrices for LLaVA and GPT-4V on Dollar Street images
      4. iv.
        Confusion matrices for LLaVA and GPT-4V on MaRVL images
    2. (b)
      Task 2 - Artifacts
      1. i.
        Number of artifacts identified using GPT-4V
A.1 Dataset details